OpenStudy (anonymous):

((x+7)/(x^(3)(10-x)))(x(x-10)/11) Simplify.

OpenStudy (anonymous):

not as bad as it looks. cancel and x top and bottom and note that \[\frac{x-10}{10-x}=-1\] so that goes too leaving you with \[-\frac{x+7}{11x^2}\]